OpenStudy (anonymous):

Algebra 1 portfolio? I'm having a hard time with it :( can someone help me?

OpenStudy (anonymous):

What exactly is a Math Portfolio?

OpenStudy (anonymous):

it's a way of doing a larger assignment in Connections Academy

OpenStudy (ryan123345):

:DD

OpenStudy (ryan123345):

sure :D

OpenStudy (anonymous):

koolio!

OpenStudy (ryan123345):

alright :3

OpenStudy (anonymous):

are you connections academy?

OpenStudy (ryan123345):

im not sure what that is lol but if u need help with algebra ill be happy to help u :)

OpenStudy (anonymous):

computer capacity is measured in bits and bytes. A bit is the smallest unit, which is a 1 or a 0, in the computers memory. A byte is 2 to the 3rd bower bits. A megabyte is 2 to the 20th power bytes. how many bits are in a megabyte?

OpenStudy (anonymous):

sorry bower is suppose to be power. lol

hero (hero):

1 byte = 8 bits 1 megabyte = 2^20 bytes = 2^20 × 8 bits

OpenStudy (anonymous):

This is awesome! thanks :)

OpenStudy (anonymous):

A gigabyte is 2^30 bytes. how many megabytes are in a gigabyte? how many bits are in a giabyte?